OpenStudy (anonymous):

help me solve this equation x/2 - y/3=1

OpenStudy (anonymous):

What should u want me to do?

OpenStudy (anonymous):

i need help with this equation

OpenStudy (anonymous):

Sure, but what i have to do with line equation?

OpenStudy (anonymous):

first of all you are going to have 2 values one for x and y value

OpenStudy (anonymous):

so you just take it apart piece by piece solve by one variable then by the next

OpenStudy (anonymous):

so do i find a common deminator

OpenStudy (anonymous):

well do you know that you have to have to values x and y?

OpenStudy (anonymous):

not that i told you

OpenStudy (anonymous):

because first off we must solve for one variable first whichever one (ill choose x) x/2 - y/3=1 x/2 - y/3 + (y/3) += 1 + (y/3)

OpenStudy (anonymous):

Wait, What are trying to explain?????

OpenStudy (anonymous):

i am trying to get a start at the equation am i doing something wrong? plz tell me

OpenStudy (anonymous):

so i may correct myself

OpenStudy (anonymous):

to continue...

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

shonie5 are you getting it?

OpenStudy (anonymous):

@iforgot thanks i have x/2 = 1 + y/2, now i'm going to simplify for x

OpenStudy (anonymous):